ABCD is a square; forces of 1 kg wt, 6 kg wt and 9 kg wt act in the directions AB, AC and AD respectively; find the magnitude of their resultant correct to the two decimal points.

Grade:11

1 Answers

Shailendra Kumar Sharma
6 years ago
Assume A be origin, AB be X axis i.e i be the unit vector and AD be y axis i.e  j  be the unit vector
so AC will be i+j/sqrt{2} 

SO net force will be {1 i +6 ((i+j)/\sqrt{2} )+ 9j  }Xg
{(1+6/\sqrt{2})i +(1+9/\sqrt{2})j}X 9.8

So magnitude 51.38 N